OpenStudy (anonymous):

(10, 11) m = -9 Would it be y=-90x+11

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (alexandervonhumboldt2):

first y=ax+b then y=-9x+b substitue point to find b: 11=-9*10+b. b=101. the equation is y=-9x+101 i explained you thousand of times that

OpenStudy (anonymous):

can someone help me on my problem

OpenStudy (alexandervonhumboldt2):

y=ax+b where a is slope and b is y-intercept

OpenStudy (anonymous):

My bad! Lol thank you. This is confusing to me.